OpenStudy (anonymous):

Evaluate -4(a+b) - 10a/b when a=8 and b=5 a. -15 b. 36 c. -53 d. 625/8 e. -68

OpenStudy (anonymous):

so what is the correct answer for 9 widgets?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

so is that the answer

OpenStudy (anonymous):

Robert why dont you explain how you got that answer so she could understand how you got the answer..

OpenStudy (anonymous):

-68 is the correct ans.

OpenStudy (anonymous):

Thanank you Sue101, but I understand now. :) thanks Skittles.

OpenStudy (anonymous):

-4(13)-10(8/3) -52-16=-68

OpenStudy (anonymous):

hey pkjh3105 answer this for me please A rectangular display case houses a square pyramid. Both have the same square base measuring x inches on a side. The top of the case is 2 inches above the tip of the pyramid which has a height of h inches. Which expression represents the volume of the empty space inside the case? Answer: a. x^3 - x^2(h-2) b. x^2-1/3x^2h c. 1/3x2^h d. 2/3x^2h e. x^2(h+2) - 1/3x^2h

OpenStudy (anonymous):

hey..hv already answrd it...its c...

OpenStudy (anonymous):

that says it is incorrect it cant be c

OpenStudy (anonymous):

In a coordinate system, points A(-2, 3) and B(6, -1) have a midpoint of X. What is the midpoint between A and X? a. (0,0) b. (2, 1) c. (0,2) d. (4,2) e. (-1,1.5

OpenStudy (anonymous):

How about this one: Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. $18-2 calc. $66- 4 calc $450- 10 calc. what is the total cost of producing 8 calc?

OpenStudy (anonymous):

b..

OpenStudy (anonymous):

A trapezoid has a base of x and height of 3 units. If there is also a square with sides measuring x , what formula would be used to calculate the total area of both figures? a. x^2 + 3x b. 4x + x^3 c. x^2 + 9x d. 9x e. x^2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

pkjha3105 If x= y-3 and y=z^2, what is x in terms of z?

OpenStudy (anonymous):

a. z^-6 b. z^6 c. z^-2/3 d. z^5 e. z^3/2

OpenStudy (anonymous):

robert